January 9, 2011

OBSERVERS: Jill Bechtold, David Sand, Marco Pedani, Michael Rutkowski, Russell Ryan

OPERATOR: David Gonzales Huerta

Afternoon: Clouds.

3:43 pm. Turned off lights in the dome.

When we started, LBC housekeeping had false error flags, and we had to restart CMU. This only took 5 minutes.

3:53 pm. Started taking biases, 5bias_bino.ob

4:06pm. Can't get RB_Science started. Calling Olga. Later figured out that RB_Science doesn't work on bias images.

4:26pm. Bias frames do not have WCS so we can't open them in ds9 with "Open-Other-Mosaic"

Filled dewars.

5:03pm. Started 25binocular_biases.ob at 12:03 UTC

5:49pm. Opening. Going to try Lucifer first. Clouds all around, but thinning. Off to point near sdss0405.

5:55pm. Executed sdss0405.Hacq.txt. Too bright out to see guide star, so we'll wait.

6:20pm Executed sdss0405.Hacq.txt. Guide star now ok.

Seeing is 5". ERGH... primary 1 deg warmer than ambient.

clouds around.

6:36. Going to point near sn09ig, and see what's happening there.

Really bad seeing there (2-3"), and clouds are getting thicker.

6:47pm. Decided to switch to LBC and take art shots, with the hope that by the time the science fields are up later in the night, the seeing will have settled down. Started transformation to LBC.

7:17pm Blue side primary mirror losing pressure. Calling for help.

7:30pm Pressure problem has been fixed on mirrors.

7:32pm Connecting LBC. Moving to M77. Art shot to see what seeing is.

7:36pm Running FPIA. 1 degree north of M77. FPIA says 2.4 arcsec seeing.

7:43pm Running M77. V and Y, 1 sec each.

7:48 V and Y in M77. 100 sec. 2.9 arcsec seeing.

7:57 SN_2009ig exposures for Milne. Bessel B and R. 240 sec per image. Airmass 1.2; Elevations 55 deg. 2.9 arcsec seeing in blue; 2.2 in red.

8:21 running FPIA so we can get second OB at this position. Claimed blue seeing is 1.8 arcsec.

8:25 SN_2009ig V and I Bessel. I is 180 sec. V is 180. V is 2.2 arcsec; I is 1.5 arcsec.

8:43 Closed for high humidity.

Red CCD controller failed, so we lost 3 I-images at end of last OB.

9:06pm Recycling power to restore Red side.

9:15pm Dewpoint better, waiting 15 minutes to see if it lasts before reopening.

9:22pm. Marco is turning the camera on and off again.

9:34pm. Dewpoint alarm again. Cannot open.

9:48pm. Turned building, wind rising.

10pm. Humidity 102%. Doesn't look good for tonight.

4:55 am. We are still closed to wind. Packing up and calling it a night.
I Attachment Action Size Date Who Comment
20110110.fwhmfwhm 20110110.fwhm manage 1 K 10 Jan 2011 - 23:12 XiaohuiFan LBC Image Quality Statistics from Jan 9, 2010 (UT Jan 10)
Topic revision: r6 - 11 Jan 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback